(y+z)p+(z+x)q =x+y


1
Expert's answer
2021-02-24T07:37:41-0500

"(y+z)p+(z+x)q=x+y"


This is Lagrange's linear equation "Pp+Qq=R"


∴ The auxiliary equation is


"\dfrac{dx}{y+z}=\dfrac{dy}{z+x}=\dfrac{dz}{x+y}"


Taking the fraction as:-


"\dfrac{dx+dy+dz}{2(x+y+z)}=\dfrac{dx-dy}{y-x}=\dfrac{dy-dz}{z-y}"


Taking first two terms-

 

"\dfrac{dx+dy+dz}{2(x+y+z)}=\dfrac{dx-dy}{y-x}"


"\dfrac{dx+dy+dz}{2(x+y+z)}=\dfrac{-d(x-y)}{x-y}"


Integrating Both side and we get,

 

"\dfrac{1}{2}log(x+y+z)=-log(x-y)+C"


Now taking last two terms-

"\dfrac{dx-dy}{y-x}=\dfrac{dy-dz}{z-y}"


"\dfrac{d(x-y)}{y-x}=\dfrac{d(y-z)}{z-y}"


Integrating Both sides and we get-

 

 "log(x-y)=log(y-2)+logc_2"


"\dfrac{x-y}{y-z}=c_2"


Hence General equation of the differential equation is-


"\phi[(x+y+z)(x-y)^2,\dfrac{(x-y)}{(y-z)}]=0"